# Adaptive Layer Splitting for Wireless LLM Inference in Edge Computing: A Model-Based Reinforcement Learning Approach

### v3

@@ V-A Experimental Setup
To reduce the computational cost of evaluating the LLM’s performance at each step, by collecting $9,718$ pieces of practical records, we derive a reward surrogate model as in Section IV-C. Our result shows that a multi-layer perceptron (MLP) yields a test loss of $0.00548$ in MSE and $0.050$ in mean absolute error (MAE), thus providing sufficient accuracy. Therefore, we use this MLP-based reward surrogate model to accelerate the evaluation process.
@@ V-B Experiment Results
We first present the performance of PPO with and without the reward surrogate model, and compare them with the baseline RL schemes (i.e., A2C [58] and DQN [59]). Fig. 5 presents the corresponding results. It can be observed from Fig. 5 that for Case H and Case A, PPO yields significantly superior performance than A2C and DQN; while for Case L, all RL approaches lead to similar performance. Besides, PPO trained with reward surrogate models closely resemble that with actual rewards. Furthermore, Table IV compares PPO with and without reward surrogate model under Case A in terms of reward, training duration, and computational resource consumption. The results indicate that the reward surrogate model significantly reduces the training time and computational resource consumption while achieving comparable rewards.
@@
Fig. 7 illustrates $500$ splitting points determined by the trained PPO agent under Case A, and complements a locally estimated scatterplot smoothing (LOESS) [60] trend line, whose slope and R-squared value provide quantitative insights into the relationship between channel conditions and splitting point decisions. It can be observed from Fig. 7 that as noise intensity increases, the agent prefers to place the splitting point further from the input layers. This strategic adjustment helps mitigate the adverse effects of noise on the model’s performance by leveraging the more robust processing capabilities of the cloud.
@@ VI Conclusion and Future Works
In this paper, we have presented an MBRL framework for dynamically optimizing the splitting point of LLMs deployed across UE and the edge, so as to enhance the efficiency and performance of LLMs under wireless network conditions. In particular, we have formulated the problem as an MDP, and introduced a reward surrogate model to significantly shorten overall training time. The experimental results have demonstrated the framework’s efficacy in managing the trade-off between inference performance and computational load at the UE. Meanwhile, comprehensive validations in mainstream open-source LLMs have clearly demonstrated that an earlier model splitting could worsen the point inference performance, which might provide an independent interest to the community.
Despite these achievements, several limitations and challenges remain. Though the validation of the impact of splitting points on the performance of some widely adopted LLMs, given the versatility of LLMs, the generality issue still awakens further attention. The lack of a more accurate channel model and the absence of communication-efficient distribution learning approaches (e.g., quantization) in data transmission also demands future research. Additionally, the scalability of our framework in larger, more complex network environments and its generalization across different LLM architectures warrant further investigation. We will explore these important directions in the future.

### v4

@@ V-A Experimental Setup
To reduce the computational cost of evaluating the LLM’s performance at each step, by collecting $9,718$ pieces of practical records, we derive a reward surrogate model as in Section IV-C. Our result shows that an MLP yields a test loss of $0.00548$ in MSE and $0.050$ in mean absolute error (MAE), thus providing sufficient accuracy. Therefore, we use this MLP-based reward surrogate model to accelerate the evaluation process.
@@ V-B Experiment Results
We first present the performance of PPO with and without the reward surrogate model and compare them with baseline RL schemes (i.e., A2C [66] and DQN [67]). Fig. 5 presents the corresponding results. It can be observed from Fig. 5 that for Case H and Case A, PPO yields significantly superior performance than A2C and DQN; while for Case L, all RL approaches lead to similar performance. Besides, PPO trained with reward surrogate models closely resemble that with actual rewards. Furthermore, Table IV compares PPO with and without reward surrogate model under Case A in terms of reward, training duration, and computational resource consumption. The results indicate that the reward surrogate model significantly reduces the training time and computational resource consumption while achieving comparable rewards.
@@
Figure 8: Comparison of training performances across different action spaces (Action Space = 3, 5, 7) for varying movement ranges.
@@
Fig. 7 illustrates $500$ splitting points determined by the trained PPO agent under Case A, and complements a locally estimated scatterplot smoothing (LOESS) [68] trend line, whose slope and R-squared value provide quantitative insights into the relationship between channel conditions and splitting point decisions. It can be observed from Fig. 7 that as noise intensity $\sigma$ increases, the agent prefers to place the splitting point further from the input layers. This strategic adjustment helps mitigate the adverse effects of noise on model performance by leveraging the cloud’s more robust processing capabilities.
In addition to the baseline experiment where the action space is limited to single-layer adjustments, we conduct further experiments with enlarged action space to evaluate the impact of larger adjustments on the training process and final rewards. As shown in Fig. 8, allowing larger adjustments (e.g., moving by ±$\textpm 2$, ±$\textpm 3$ layers) leads to slower convergence but ultimately achieves higher rewards. This trade-off suggests that while the larger action space can explore a wider range of configurations, they may require more training steps to stabilize. However, in both experimental and practical scenarios, single-layer adjustments offer notable advantages. They provide fine-grained control over the splitting point, allowing the model to quickly adapt to changes in channel conditions. This is particularly beneficial in dynamic environments where frequent and subtle adjustments are necessary to maintain optimal performance.
@@ VI Conclusion and Future Works
In this paper, we have presented an MBRL framework for dynamically optimizing the splitting point of LLMs deployed across UE and the edge, so as to enhance the efficiency and performance of LLMs under wireless network conditions. In particular, we have formulated the problem as an MDP, and introduced a reward surrogate model to significantly shorten overall training time. The experimental results have demonstrated the framework’s efficacy in managing the trade-off between inference performance and computational load at the UE. Meanwhile, comprehensive validations in mainstream open-source LLMs have clearly demonstrated that an earlier model splitting could worsen the point inference performance, which might provide an independent interest to the community. Our proposed framework offers a structured approach to dynamically deploying LLMs across heterogeneous devices. In practical applications, such as in smart cities and industrial IoT, this framework can enhance the flexibility of LLM deployment while alleviating the computational constraints associated with running LLMs on edge devices.
Despite these achievements, several limitations and challenges remain. Though the validation of the impact of splitting points on the performance of some widely adopted LLMs, given the versatility of LLMs, the generality issue still awakens further attention. The lack of a more accurate channel model and the absence of communication-efficient distribution learning approaches (e.g., quantization) in data transmission also demands future research. Also, future research could explore adaptive mechanisms that dynamically adjust the action space based on the current learning phase or environmental conditions, thereby balancing the need for quick convergence and high reward attainment. Additionally, further investigation is required into the scalability of our framework in larger, more complex network environments and its generalization across different LLM architectures. We will explore these important directions in the future.